Graph each function. Then identify the vertex, two points, domain, and range.


f(x) = |-1/5(x + 5)| + 2


1
Expert's answer
2021-10-06T12:17:23-0400

Solution:

The graph of "f(x)=|-\frac{\left(x\:+\:5\right)}{5}|\:+\:2" is



From the graph:

"\mathrm{Domain\:of\:}\:\frac{\left|x+5\right|}{5}+2\::\quad \begin{bmatrix}\mathrm{Solution:}\:&\:-\infty \:<x<\infty \\ \:\mathrm{Interval\:Notation:}&\:\left(-\infty \:,\:\infty \:\right)\end{bmatrix}"

"\mathrm{Range\:of\:}\frac{\left|x+5\right|}{5}+2:\quad \begin{bmatrix}\mathrm{Solution:}\:&\:f\left(x\right)\ge \:2\:\\ \:\mathrm{Interval\:Notation:}&\:[2,\:\infty \:)\end{bmatrix}"

Vertex ": (-5,2)"

Two points ":(0,3),(5,4)"